I need to solve this by losing the fraction
2n-3 over 2 = 3/4

Answer by mananth(16949)

2n-3 / 2= 3/4
eliminate the denominator on the left hand side of the equation by multiplying by 2
2n-3 / 2 * 2 = 3/4 * 2
2n-3 = 3/2
similarly for right hand side of the equation multiply by 2
2(2n-3) = 3/2 * 2
4n-6=3
4n=6+3
4n=9
n=9/4
---------------
2n-3 / 2= 3/4
OR later you could just do it this way

4(2n-3) = 2*3
8n-12 = 6
8n=6+12
8n=18
n=18/8
n= 9/4